# LapSum – One Method to Differentiate Them All: Ranking, Sorting and Top-k Selection

### v1

@@ 3.4 Soft permutations
The task in the problem of soft permutation is to generalize the hard permutation matrix into a differentiable double stochastic matrix.
Recall that the permutation matrix corresponding to a permutation $(s_{i})_{i=1..n}$ of the indices $(1,\ldots,n)$ is a square binary matrix that represents the reordering of elements in the sequence. More precisely, the permutation matrix $P$ is a matrix $n\times{}n$ where:
    each row and column contains exactly one entry of $1$, and all other entries are 00,
@@
More precisely, to construct $P$ first create a zero $n\times n$ matrix, and then for each $i\in\{1,2,\dots,n\}$, place an $1$ in the $(i,p_{i})$-th position.
@@
By $\mathrm{B}_{n}$ we denote the space of all doubly stochastic matrices of size $n\times n$, where we recall that a doubly stochastic matrix is a square matrix of non-negative real numbers where each row and each column sum to 1.
@@
where $b_{-1}=-\infty$, $b_{l}=F^{-1}(l)$ for $l=0..n-1$, $b_{n}=\infty$.
Visualization of the definition is given in Fig. 4. We proceed to theorem which gives justification of the defined function.
@@
${F\text{-}\mathrm{Perm}}_{\alpha}(r)\to\mathrm{permutation\_matrix}(r)\text{ as }\alpha\to 0^{+}.$
@@
We prove that ${F\text{-}\mathrm{Perm}}_{\alpha}$ is double stochastic. Fix an arbitrary $j\in\{0,\ldots,n-1\}$. Then
$\sum_{i=0}^{n-1}(F_{\alpha}(b_{i+1}-r_{j})-F_{\alpha}(b_{i}-r_{j}))=F_{\alpha}(\infty)-F_{\alpha}(-\infty)=1.$
@@
$\sum_{i=0}^{n-1}(F_{\alpha}(b_{i+1}-r_{j})-F_{\alpha}(b_{i}-r_{j}))=$
$={F\text{-}\mathrm{Sum}}(b_{j+1};r)-{F\text{-}\mathrm{Sum}}(b_{j};r)=(j+1)-j=1.$
The proof of the limiting case $\alpha\to 0^{+}$ can be done similarly to the reasoning from Theorem B.1. ∎
@@ 4 Theory behind $\mathrm{LapSum}$
In the previous section we have shown that we can construct a general soft-order theory based on an arbitrary cumulative density function $F$. However, such an approach is in practice inapplicable, since the complexity of evaluating ${F\text{-}\mathrm{Sum}}(x;r)$ in $n$ points for sequence $r$ of size $n$ is in general quadratic. Moreover, to efficiently apply the theory, we need to be able to compute the inverse function ${F\text{-}\mathrm{Sum}}^{-1}(w;r)$, which would usually need using some nontrivial iterative procedure. Finally, to use the model in neural networks, we would optimally need the derivatives computed in complexity $O(n\log{}n)$.
@@ 4.1 Where the magic happens …
${Lap\text{-}\mathrm{Sum}}_{\alpha}(x;r)=\sum_{i=0}^{n-1}Lap_{\alpha}(x-r_{i}),$
@@
The function ${Lap\text{-}\mathrm{Sum}}$ will enable to solve all classical soft order problems – we shall denote the family of methods constructed with the use of ${Lap\text{-}\mathrm{Sum}}$ function as $\mathrm{LapSum}$. We show two crucial facts which enable the efficient use of $\mathrm{LapSum}$ in soft order problems:
  • we can evaluate ${Lap\text{-}\mathrm{Sum}}$ in $n$ points in complexity $O(n\log n)$.
@@
Suppose that $r=(r_{i})_{i=0..n-1}\subset\mathbb{R}$ is an increasing sequence. We can compute the values of ${Lap\text{-}\mathrm{Sum}}$ for an increasing sequence $X=(x_{j})_{j=0..m-1}\subset$ with a complexity of $O(n+m)$.
To prove this theorem we will need following proposition.
@@
[…]
@@
$\begin{array}[]{l}{Lap\text{-}\mathrm{Sum}}(x)=\\[4.30554pt] =\tfrac{1}{2}a_{k}\exp(\tfrac{x-r_{k+1}}{\alpha})-\tfrac{1}{2}b_{k+1}\exp(\tfrac{r_{k}-x}{\alpha})+c_{k+1}\\[4.30554pt] \hfill\text{ for }x\in[r_{k},r_{k+1}],k=-1..(n-1).\end{array}$ (4)
where sequences $(a_{i})_{i=-1..n-1},(b_{i})_{i=0..n}\subset\mathbb{R}$ and $(c_{i})_{i=0..n}\subset\mathbb{N}$ are defined iteratively by the formulae:
  • $a_{N-1}=0$, $a_{k-1}=\exp(\frac{r_{k-1}-r_{k}}{\alpha})\cdot(1+a_{k})$,
  • $b_{0}=0$, $b_{k+1}=\exp(\frac{r_{k}-r_{k+1}}{\alpha})\cdot(1+b_{k})$,
  • $c_{0}=0$, $c_{k+1}=1+c_{k}$.
@@
Given that $F$ on the intervals $(-\infty,0]$ and $[0,\infty)$ is composed of constants, $\exp(x/\alpha)$, and $\exp(-x/\alpha)$, this property extends to function ${Lap\text{-}\mathrm{Sum}}_{\alpha}(x;r)$ for every interval $[r_{k},r_{k+1}]$. Consequently, there exist sequences $a_{k},b_{k},c_{k}$ that satisfy the formula given by (4). We aim to demonstrate that the proposition’s recurrent formulas are valid.
Consider the interval $(-\infty,r_{0})$. Clearly, $c_{0}$ is zero because the expression for ${Lap\text{-}\mathrm{Sum}}$ in this range is composed of exponentials. Assuming that the expression for $c_{k}$ holds, it follows that in the next interval, we increase by $1$. This logic is uniformly applied to all subsequent $c_{k}$. Now, consider the sequence $a_{k}$. Initially, note that $a_{n-1}=0$. Using a similar reasoning, we derive the entire recursive formula. ∎
Now we are ready to present the proof of Theorem 4.1.
@@
Since the sequence $(r_{i})_{i=0..n-1}$ is ordered, by the previous proposition we can compute the sequences $a_{i},b_{i},c_{i}$ in complexity $O(n)$. Now in complexity $O(k)$ we can find $n_{l}$ for $l=0..k-1$ such that $x_{l}\in[r_{n_{l}-1},r_{n_{l}}]$. Finally by the previous proposition we get that ${Lap\text{-}\mathrm{Sum}}_{\alpha}(x_{l})$ is given by
$\tfrac{1}{2}a_{k}\exp(\tfrac{x_{l}-r_{k+1}}{\alpha})-\tfrac{1}{2}b_{k+1}\exp(\tfrac{r_{k}-x_{l}}{\alpha})+c_{k+1},$
which directly implies that we can compute all this values in $O(k)$ complexity. ∎
@@
We now show that function $(0,n)\ni w\to{Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w;r)$ can be computed with a simple straightforward formula. For clarity, we present the formulae for $\alpha=1$.
Compute first $w_{i}={Lap\text{-}\mathrm{Sum}}{(r_{i})}$ for all $i=0..n-1$. If $w\in(-\infty,w_{0}]$, then as on this interval ${Lap\text{-}\mathrm{Sum}}(x)=\tfrac{a_{0}}{2}\exp(x-r_{0})$, solution to ${Lap\text{-}\mathrm{Sum}}(x)=w$ is given by
@@
If $w\in[r_{i-1},r_{i}]$ for some $0<i<n$, since
${Lap\text{-}\mathrm{Sum}}(x)=\tfrac{1}{2}$ $a_{k}\exp(x-r_{k+1})$
-- $\tfrac{1}{2}b_{k+1}\exp(r_{k}-x)+c_{k+1},$
by direct calculations, the solution to ${Lap\text{-}\mathrm{Sum}}(x)=w$ is
$\begin{array}[]{l}x=r_{k+1}-\log a_{k}+\\ +\log(w-c_{k+1}+\sqrt{|w-c_{k+1}|^{2}+a_{k}b_{k+1}\cdot e^{r_{k}-r_{k+1}}}).\end{array}$
If $w\in[r_{n},\infty)$, we analogously compute that
$x=r_{n-1}-\log 2-\log(c_{n}-w)+\log b_{n}.$
@@ 4.2 Efficient computation of derivatives
Let us start with the case of top-w selection, namely for the function ${Lap\text{-}\mathrm{Top}}_{\alpha}(r;w)$. We define the function $P:\mathbb{R}^{n}\ni r=(r_{i})\to p=(p_{i})\in\Delta_{w}\subset$, where
$p_{i}=Lap_{\alpha}(b-r_{i})\text{ where }b={Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w;r).$
One can show (see Appendix Sec. C.2), that the derivative of $p$ with respect to $r$ is given by the matrix
@@
$q=\mathrm{softmax}(-|\tfrac{b-r_{0}}{\alpha}|,\ldots,-|\tfrac{b-r_{N-1}}{\alpha}|).$
For optimizing purposes, utilizing such a constructed derivative would result in a complexity of at least $O(n^{2})$. However, note that to apply gradient optimization it is actually enough to compute $v^{T}D$ and $Dv$ for an arbitrary vector $v$. By executing direct calculations, one can trivially derive from (5) the formulae
@@
Another important function, which is necessary to calculate soft-sorting and permutations, is the function (for fixed $k\leq n$)
$\begin{array}[]{l}L(w_{0},\ldots,w_{k-1};r)=\\ =({Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w_{0};r),\ldots,{Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w_{k-1};r)).\end{array}$
In this instance, calculating the derivative of $L$ efficiently, in particular the left and right multiplication of the derivative by row and column vectors, respectively, can be also implemented within a time complexity of $O(n\log{}n)$.
@@ 5 Experiments and results
The main incentive of this paper was to design a new easy to implement method given by closed-form, which would constitute a tool for use in all the order tasks. We wanted $\mathrm{LapSum}$ to outperform other methods in terms of time and memory complexity, while obtaining the general SOTA in the standard experiments.
@@ 5.1 Experiments on soft-top-k methods
Our evaluation consists of training from scratch on CIFAR-100 and fine-tuning on ImageNet-1K and ImageNet-21K-P datasets. These evaluations aim to assess $\mathrm{LapSum}$’s influence of top-1 and top-5 training on accuracy metrics. The experiment highlights the effect of the smaller number of classes in CIFAR-100 compared to the larger count of 1000 in ImageNet-1K and an even greater count of 10450 in ImageNet-21K-P. We employ $P_{K}$ settings akin to those in Petersen et al. (2022b). Experiment details are available in Appendix D.
Table 1: Performance comparison of permutation-based methods on CIFAR-100 using ResNet18. The table shows ACC@1 (standard accuracy) and ACC@5 (top-5 accuracy). Bold and italic values denote the best and second-best results, respectively. $P_{K}$ represents the training probability distribution. Methods marked with (*) are based on Berrada et al. (2018).
Method $P_{K}$ CIFAR-100
@@
Lap-Top-k (ours) [.2,.2,.2,.2,.2] 63.22 $\boldsymbol{87.10}$
Refer to Tab. 1 for the CIFAR-100 results, where we used $P_{K}=[0.,0.,0.,0.,1.]$ for top-5 learning. Furthermore, we used [.2,.2,.2,.2,.2].2.2.2.2.2[.2,.2,.2,.2,.2][ .2 , .2 , .2 , .2 , .2 ] to have a direct comparison with the results given in Petersen et al. (2022b). Our experiments demonstrated superior or at least comparable results. The results are consistently stable and are fast to obtain.
Table 2: Results on ImageNet-1K and ImageNet-21K-P for fine-tuning the head of ResNeXt-101 (Mahajan et al., 2018). We report ACC@1 and ACC@5 accuracy metrics, averaged over 10 seeds for ImageNet-1K and 2 seeds for ImageNet-21K-P. Methods are evaluated using different $P_{K}$ configurations, highlighting the effectiveness in optimizing ranking-based losses. Best performances are in bold, with second best in italics. Methods labeled with (*) are based on Berrada et al. (2018).
Method $P_{K}$ ImgNet-1K ImgNet-21K-P
@@
Lap-Top-k (our) [.5,0,0,0,.5] 86.25 97.93 40.48 71.05
Figure 5: Upper: the relationship between the data dimension $n$ (horizontal axis) and the maximum memory usage and computation time (vertical axis), with $k=n/2$ in this figure. Computations for forward/backward process shown here were performed on a CPU (see Appendix Figs. 11, 13, 10 and 12 for more). Below: the critical confidence diagram showing the statistical confidence of the above results (Demšar, 2006). $\mathrm{LapSum}$ approach is comparable to the best few (better on the right, statistically comparable joined with horizontal segments).
Tab. 2 summarizes the outcomes for fine-tuning on ImageNet-1K and ImageNet-21K-P, similarly to the setup in Petersen et al. (2022b). The ImageNet-21K-P results with  $\mathrm{LapSum}$ surpass alternative methods. $\mathrm{LapSum}$ yields slightly superior results faster due to $\mathrm{LapSum}$’s efficient time and memory management in handling this intricate task. The results of ImageNet-1K are on par with those of other methods. This shows that $\mathrm{LapSum}$is particularly well suited for high-dimensional tasks, being fast and easy to use for other tasks. At the same time, $\mathrm{LapSum}$ has a differentiable parameter $\alpha$ (see Sec. 3.1), which can be used to fine-tune individual solutions. The results of these experiments are presented in Appendix D.
Petersen et al. (2022b) suggests the algorithm a predetermined number $m$ of top scores for differential ranking for enhancing its effectiveness. In contrast, our investigation shows that $\mathrm{LapSum}$ operates effectively without the requirement of this proposal. It excels with any dimension $n$.
@@
We have also performed experiments on the accuracy of direct approximations of $k$ value using $\mathrm{LapSum}$, see Fig. 7. Our proposition consistently gives better approximations than competing approaches, irrespective of a given problem dimension. Fig. 6 shows the impact of $\mathrm{LapSum}$ $\alpha$ trainable parameter value on training on the resulting in harder or softer solutions (see Fig. 8).
@@
The protocol followed was adapted from (Xie et al., 2020), we use the code kindly made available by authors are copy their results as baselines. Tab. 3 shows test accuracies on MNIST and CIFAR-10. We see that Lap-top-k does not deteriorate the model performance. We provide additional details in Appendix D.
Table 3: Accuracy of $k$NN-based methods on MNIST and CIFAR-10 comparing standard $k$NN with enhanced variants. Bold and italic values stand for best and second-best results. Our proposed $k$NN+Lap-Top-$k$ shows strong performance, especially on CIFAR-10, showing the benefits of Laplacian-based top-k optimization.
@@
$k$NN 97.2 35.4
$k$NN+PCA 97.6 40.9
$k$NN+AE 97.6 44.2
$k$NN+pretrained CNN 98.4 91.1
@@
$k$NN+NeuralSort 99.5 90.7
$k$NN+OT 99.0 84.8
$k$NN+Softmax $k$ times 99.3 92.2
@@
$k$NN+SOFT Top-$k$ 99.4 92.6
$k$NN+Lap-Top-$k$ (ours) 99.4 92.2
[…]

### v2

@@ 3.4 Soft permutations
The task of the soft permutation problem is to generalize the hard permutation matrix to a differentiable double stochastic matrix.
Recall that the permutation matrix corresponding to a permutation $(s_{i})_{i=1..n}$ of the indices $(1,\ldots,n)$ is a square binary matrix that represents the reordering of elements in the sequence. Precisely, the permutation matrix $P$ is a matrix $n\times{}n$ where:
    each row and column contains exactly one entry of $1$, and all other entries are 0,
@@
To construct $P$, first create a zero matrix $n\times n$, and then, for each $i\in\{1,2,\dots,n\}$, place a $1$ in the $(i,p_{i})$-th position.
@@
By $\mathrm{B}_{n}$ we denote the space of all doubly stochastic matrices of size $n\times n$, where we recall that a doubly stochastic matrix is a square matrix of non-negative real numbers where each row and each column sum up to 1.
@@
where $b_{-1}=$-$\infty$, $b_{i}=F^{-1}(i)$ for $i=0,\dots,n-1$, $b_{n}=\infty$.
Visualization of the definition is given in Fig. 5. We proceed to theorem, which gives justification of the defined function.
@@
${F\text{-}\mathrm{Perm}}_{\alpha}(r)\to\mathit{permutation\_matrix}(r)\text{ as }\alpha\to 0^{+}.$
@@
We prove that ${F\text{-}\mathrm{Perm}}_{\alpha}(r)$ is doubly stochastic. Let an arbitrary $j\in\{0,\ldots,n-1\}$. Then
$\sum_{i=0}^{n-1}\left(F_{\alpha}(b_{i+1}-r_{j})-F_{\alpha}(b_{i}-r_{j})\right)$
$F_{\alpha}(\infty)-F_{\alpha}(-\infty)=1.$
@@
$\sum_{j=0}^{n-1}\left(F_{\alpha}(b_{i+1}-r_{j})-F_{\alpha}(b_{i}-r_{j})\right)=$
${\;\;}={F\text{-}\mathrm{Sum}}(r,b_{j+1})-{F\text{-}\mathrm{Sum}}(r,b_{j})=(j+1)-j=1.$
The proof of the limiting case $\alpha\to 0^{+}$ can be done similarly to the reasoning from Theorem B.1 in the Appendix. ∎
@@ 4 Theory behind $\mathrm{LapSum}$
In the previous section we have constructed a general soft-order theory based on an arbitrary cumulative density function $F$. However, such an approach is in practice unapplicable, since the complexity of evaluating ${F\text{-}\mathrm{Sum}}(r,x)$ for a sequence $r$ of size $n$ is in general quadratic. Moreover, to efficiently apply the theory, we need to be able to compute the inverse function ${F\text{-}\mathrm{Sum}}^{-1}(r,w)$. Finally, to use the model for neural networks, we would optimally need the derivatives computed in complexity $O(n\log{}n)$.
@@ 4.1 Where the magic happens …
${Lap\text{-}\mathrm{Sum}}_{\alpha}(r,x)=\sum_{i=0}^{n-1}Lap_{\alpha}(x-r_{i}),$
@@
The function $Lap\text{-}\mathrm{Sum}$} will be crucial in solving all previously mentioned soft-order problems – we shall denote the family of methods constructed with the use of $Lap\text{-}\mathrm{Sum}$} function as $\mathrm{LapSum}$. We show two crucial facts which enable the efficient use of $\mathrm{LapSum}$ in soft-order problems:
  • we can evaluate $Lap\text{-}\mathrm{Sum}$} at $n$ points in complexity $O(n\log n)$,
  • the computation of $Lap\text{-}\mathrm{Sum}$} can be parallelized,
@@
Suppose that $r=(r_{i})_{i=0..n-1}\subset\mathbb{R}$ is an increasing sequence. We can compute the values of $Lap\text{-}\mathrm{Sum}$} for an increasing sequence $x=(x_{j})_{j=0..m-1}\subset\mathbb{R}$ with a complexity of $O(n+m)$.
To prove this theorem, we will need the following proposition.
@@
Assume that $(r_{i})_{i=0..n-1}$ is an increasing sequence and $\alpha>0$, where we additionally put $r_{-1}=-\infty,r_{n}=\infty$. [^3] The case for $\alpha<0$ can be treated analogously.
@@
$\begin{array}[]{l}{Lap\text{-}\mathrm{Sum}}(x)=\\[4.30554pt] =\tfrac{1}{2}a_{j}\exp\left(\tfrac{x-r_{j+1}}{\alpha}\right)-\tfrac{1}{2}b_{j+1}\exp\left(\tfrac{r_{j}-x}{\alpha}\right)+c_{j+1}\\[4.30554pt] \hfill\text{ for }x\in[r_{j},r_{j+1}],j=-1..(n-1).\end{array}$ (5)
where sequences $(a_{i})_{i=-1..n-1},(b_{i})_{i=0..n}\subset\mathbb{R}$ and $(c_{i})_{i=0..n}\subset\mathbb{N}$ are defined iteratively by the formulae
  • $a_{N-1}=0$, $a_{k-1}=\exp(\frac{r_{k-1}-r_{j}}{\alpha})\cdot(1+a_{j})$,
  • $b_{0}=0$, $b_{j+1}=\exp(\frac{r_{j}-r_{j+1}}{\alpha})\cdot(1+b_{j})$,
  • $c_{0}=0$, $c_{j+1}=1+c_{j}$.
@@
Given that $F$ on the intervals $(-\infty,0]$ and $0,\infty)$ is composed of constants, $\exp(x/\alpha)$, and $\exp(-x/\alpha)$, this property extends to function ${Lap\text{-}\mathrm{Sum}}_{\alpha}(x;r)$ for every interval $r_{j},r_{j+1}]$. Consequently, there exist sequences $a_{j},b_{j},c_{j}$ that satisfy formula (5). We aim to demonstrate that the proposition’s recurrent formulas are valid.
Consider the interval $(-\infty,r_{0})$. Clearly, $c_{0}$ is zero because the expression for $Lap\text{-}\mathrm{Sum}$} in this range is composed of exponentials. Assuming that the expression for $c_{j}$ holds, it follows that in the next interval, we increase by $1$. This logic is uniformly applied to all subsequent $c_{j}$. Now, consider the sequence $a_{j}$. Initially, note that $a_{n-1}=0$. Using a similar reasoning, we derive the entire recursive formula. ∎
We are now ready to present the proof of Theorem 4.1.
@@
Since the sequence $(r_{i})_{i=0..n-1}$ is ordered, then, according to the above proposition, we can compute the sequences $a_{i},b_{i},c_{i}$ in complexity $O(n)$. Now in complexity $O(n+m)$ we can find $n_{l}$ for $l=0,\dots,j-1$ such that $x_{l}\in[r_{n_{l}-1},r_{n_{l}}]$. Finally, thanks to the proposition, we obtain ${Lap\text{-}\mathrm{Sum}}_{\alpha}(x_{l})$ is given by
$\tfrac{1}{2}a_{j}\exp(\tfrac{x_{l}-r_{j+1}}{\alpha})-\tfrac{1}{2}b_{j+1}\exp(\tfrac{r_{j}-x_{l}}{\alpha})+c_{j+1},$
which directly implies that we can compute all these values in $O(n+m)$ complexity. ∎
Parallelization
Observe that the iterative formulas for $a_{i},b_{i}$ and $c_{i}$ can be computed in parallel by applying the standard prefix scan approach.
@@
We now show that function $(0,n)\ni w\to{Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k)$ can be computed with a simple straightforward formula. For clarity, we present the formulae for $\alpha=1$.
First compute $w_{i}={Lap\text{-}\mathrm{Sum}}{(r_{i})}$ for all i=0..n1i=0..n-1. If $k\in(-\infty,r_{0}]$, then as on this interval ${Lap\text{-}\mathrm{Sum}}(x)=\tfrac{a_{0}}{2}\exp(x-r_{0})$, solution to ${Lap\text{-}\mathrm{Sum}}(x)=k$ is given by
@@
If $k\in[r_{i-1},r_{i}]$ for some $0<i<n$ then, since
${Lap\text{-}\mathrm{Sum}}(x)=\tfrac{1}{2}a_{j}\exp(x$- $r_{j+1})$
- $\tfrac{1}{2}b_{j+1}\exp(r_{j}-x)+c_{j+1},$
by direct calculations, the solution for ${Lap\text{-}\mathrm{Sum}}(x)=k$ is
$\begin{array}[]{l}x=r_{j+1}-\log a_{j}+\\ +\log(k-c_{j+1}+\sqrt{\left|k-c_{j+1}\right|^{2}+a_{j}b_{j+1}\cdot e^{r_{j}-r_{j+1}}}).\end{array}$
If $k\in[r_{n},\infty)$, we analogously compute
$x=r_{n-1}-\log 2-\log(c_{n}-k)+\log b_{n}.$
@@ 4.2 Efficient computation of derivatives
Let us start with the case of top-k selection, namely for function ${Lap\text{-}\mathrm{Top}}_{\alpha}(r,k)$. Define the function $P:\mathbb{R}^{n}\ni r=(r_{i})\to p=(p_{i})\in\Delta_{k}\subset[0,1]^{n}$, where
$p_{i}=Lap_{\alpha}(b-r_{i})\text{ where }b={Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k).$
It can be shown (see Sec. C.2), that the derivative of $p$ with respect to $r$ is given by the matrix
@@
$q=\mathrm{softmax}(-\left|\tfrac{b-r_{0}}{\alpha}\right|,\ldots,-\left|\tfrac{b-r_{N-1}}{\alpha}\right|).$
For optimizing purposes, utilizing such a constructed derivative would result in a complexity of at least $O(n^{2})$. However, note that to apply gradient optimization it is actually enough to compute $v^{T}D$ and DvD\,v for an arbitrary vector $v$. By executing direct calculations, one can trivially derive from (6) the formulae
@@
Another important function, necessary to calculate soft-sorting and permutations is, for fixed $k{}\leq{}n$, the function
$\begin{array}[]{l}L(r,k_{0},\ldots,k_{j-1})=\\ {\;\;\;}=({Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k_{0}),\ldots,{Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k_{j-1})).\end{array}$
In this instance, calculating the derivative of $L$ efficiently, in particular the left and right multiplication of the derivative by row and column vectors, respectively, can also be implemented within a time complexity of $O(n\log{}n)$.
@@ 5 Experiments and results
The main incentive of this paper was to design a new and simple to implement closed-form method that would constitute a tool for use in all the order tasks. We wanted $\mathrm{LapSum}$ to outperform other methods in terms of time and memory complexity, while obtaining the general SOTA in the standard experiments.
@@ 5.1 Experiments on soft-top-k methods
Our evaluation consists of training from scratch on CIFAR-100 and fine-tuning on ImageNet-1K and ImageNet-21K-P datasets. These evaluations aim to assess $\mathrm{LapSum}$’s influence of top-1 and top-5 training on accuracy metrics. The experiment highlights the effect of the smaller number of classes in CIFAR-100 compared to the larger count of 1000 in ImageNet-1K and an even greater count of 10450 in ImageNet-21K-P. We employ $P_{j}$ settings akin to those in Petersen et al. (2022b). Experiment details are available in Appendix D.
Table 1: Performance comparison of permutation-based methods on CIFAR-100 using ResNet18. The table shows ACC@1 (standard accuracy) and ACC@5 (top-5 accuracy). Bold and italic values denote the best and second-best results, respectively. $P_{j}$ represents the training probability distribution. Methods marked with (*) are based on Berrada et al. (2018).
Method $P_{j}$ CIFAR-100
@@
Lap-Top-k (ours) [.2,.2,.2,.2,.2] 64.53 $\boldsymbol{88.51}$
Refer to Tab. 1 for the CIFAR-100 results, where we used $P_{j}=[0.,0.,0.,0.,1.]$ for top-5 learning. Furthermore, we used $[.2,.2,.2,.2,.2]$ to have a direct comparison with the results given in Petersen et al. (2022b). Our experiments demonstrated superior or at least comparable results. The results are consistently stable and are fast to obtain.
Table 2: Results on ImageNet-1K and ImageNet-21K-P for fine-tuning the head of ResNeXt-101 (Mahajan et al., 2018). We report ACC@1 and ACC@5 accuracy metrics, averaged over 10 seeds for ImageNet-1K and 2 seeds for ImageNet-21K-P. Methods are evaluated using different $P_{j}$ configurations, highlighting the effectiveness in optimizing ranking-based losses. Best performances are in bold, with second best in italics. Methods labeled with (*) are based on Berrada et al. (2018).
Method $P_{j}$ ImgNet-1K ImgNet-21K-P
@@
Lap-Top-k (our) [.5,0,0,0,.5] 86.28 97.93 40.48 71.05
Tab. 2 summarizes the outcomes of fine-tuning on ImageNet-1K and ImageNet-21K-P, similarly to the setup in Petersen et al. (2022b). The ImageNet-21K-P results show  $\mathrm{LapSum}$ surpass alternative methods. $\mathrm{LapSum}$ yields slightly superior results faster due to $\mathrm{LapSum}$’s efficient time and memory management in handling this intricate task. The results of ImageNet-1K are on par with those of other methods. This demonstrates that $\mathrm{LapSum}$ is particularly well suited for high-dimensional tasks, being fast and easy to use for other tasks. At the same time, $\mathrm{LapSum}$ has a differentiable parameter $\alpha$ (see Sec. 3.1), which can be used to fine-tune individual solutions. The results of these experiments are presented in Appendix D.
Petersen et al. (2022b) algorithm suggests a predetermined number of $m$ best results to compute for differential ranking. This increases its effectiveness.
@@
Fig. 6 shows the impact of $\mathrm{LapSum}$ $\alpha$ trainable parameter value on training on the resulting in harder or softer solutions (see Appendix DFig. 8). We have also performed experiments on the accuracy of direct approximations of $k$ value using $\mathrm{LapSum}$, see Fig. 7. Our proposition consistently gives better approximations than competing approaches, irrespective of a given problem dimension.
@@
The protocol followed was adapted from (Xie et al., 2020). [^4] We use the code kindly made available by authors.  Tab. 3 shows test accuracies on MNIST and CIFAR-10. We see that Lap-top-k does not deteriorate the model performance. We provide additional details in Appendix D.
Table 3: Accuracy of kkNN-based methods on MNIST and CIFAR-10 comparing standard kkNN with enhanced variants. Bold and italic values stand for best and second-best results. Our proposed kkNN+Lap-Top-$k$ shows strong performance, especially on CIFAR-10, showing the benefits of Laplacian-based top-k optimization.
@@
kkNN 97.2 35.4
kkNN+PCA 97.6 40.9
kkNN+AE 97.6 44.2
kkNN+pretrained CNN 98.4 91.1
@@
kkNN+NeuralSort 99.5 90.7
kkNN+OT 99.0 84.8
kkNN+Softmax $k$ times 99.3 92.2
@@
kkNN+SOFT Top-$k$ 99.4 92.6
kkNN+Lap-Top-$k$ (ours) 99.4 92.2
Figure 7: The error calculated as $\text{Error(k)}=|\sum_{i=1}^{n}p_{i}-k|$, where $p_{i}$, as a function of $k$ (the value defining Top-$k$ in the respective methods) and $n$ (the data dimension). The plots correspond to the errors obtained in the computations presented in Fig. 2.